You leave your $50,000 a year job and invest your savings of $40,000 (on which you were earning 10 percent annual interest) to start your own business. In the first year, your business generated $100,000 in total revenue and had to pay $40,000 in explicit costs. The economic profit generated by this gym was

a.

greater than your accounting profit because economic profit is always greater than accounting profit.

b.

$6,000

c.

$10,000

d.

$60,000
Answer #1

Answer
Accounting profit =total revenue - explicit costs
=100000-40000
=60000
Economic profit =accountig profit - implicit costs
implict costs = salary of the best alternative =the job left =$50000
Interest on the investments =40000*0.1=$4000
Total implicit costs =50000+4000=54000
Economic profit =60000-54000
=$6000
the economic profit is $6000
Option b